# A Comprehensive Evaluation of Parameter-Efficient Fine-Tuning on Method-Level Code Smell Detection

### v1

@@ Abstract.
Code smells, which are suboptimal coding practices that can potentially lead to defects or maintenance issues, can negatively impact the quality of software systems. Most existing code smell detection methods rely on heuristics-based or machine learning (ML) and deep learning (DL)-based techniques. However, these techniques have several drawbacks (e.g., unsatisfactory performance). Large Language Models (LLMs) have garnered significant attention in the software engineering (SE) field, achieving state-of-the-art performance across a wide range of SE tasks. Parameter-Efficient Fine-Tuning (PEFT) methods, which are commonly used to adapt LLMs to specific tasks with fewer parameters and reduced computational resources, have emerged as a promising approach for enhancing the performance of LLMs in various SE tasks. However, LLMs have not yet been explored for code smell detection, and their effectiveness for this task remains unclear. Furthermore, no comprehensive investigation has been conducted on the efficiency of PEFT methods for method-level code smell detection. In this regard, we systematically evaluate the effectiveness of state-of-the-art PEFT methods on both small and large Language Models (LMs) for method-level code smell detection. To begin, we constructed high-quality Java code smell datasets sourced from GitHub. We then fine-tuned four small LMs and six LLMs using various PEFT techniques, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$, for code smell detection. Our comparison against full fine-tuning revealed that PEFT methods not only achieve comparable or better effectiveness but also consume less peak GPU memory. Our analysis further explored the performance of small LMs versus LLMs in the context of code smell detection. Surprisingly, we found that LLMs did not outperform small LMs in this specific task, suggesting that smaller models may be more suited for method-level code smell detection. We also investigated the impact of varying hyper-parameter settings and the availability of training data. Our results show that increasing the number of trainable parameters does not necessarily improve the effectiveness of PEFT methods for code smell detection. In contrast, we observed significant performance gains when the training dataset size was increased. Finally, our results show that LMs combined with PEFT methods outperform traditional heuristic-based detectors, advanced DL-based approaches, and GPT-3.5-turbo in method-level code smell detection. Specifically, MCC values improve by 13.24% to 50.59% for Complex Conditional and 11.22% to 54.32% for Complex Method, demonstrating a more effective solution for identifying these code smells. In conclusion, our work provides a comprehensive evaluation of PEFT methods on both small and large LMs for method-level code smell detection. The results highlight the efficiency and effectiveness of PEFT methods in this domain, offering valuable insights for optimizing LMs for method-level code smell detection. We suggest that the choice of PEFT method for code smell detection should be guided by the model used, data availability, and computational resources. Our replication package is available at (replicationPackage, ), enabling others to reproduce and extend our findings.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0ccs: Software and its engineering Software maintenance tools
@@ 1. Introduction
To address these issues, many researchers have incorporated machine learning (ML) and deep learning (DL) algorithms to mitigate the performance instability in code smell detection that arises from the manual threshold-setting process in traditional detectors (fontana2013code, ; arcelli2016comparing, ; sharma2021code, ; liu2021deep, ). By learning to discern between smelly and non-smelly code elements, ML and DL techniques offer a solution to the subjectivity problem inherent in heuristics-based methods (nucci2018detecting, ). However, despite these advancements, ML and DL methods still face significant challenges in achieving considerable performance for code smell detection (pereira2022code, ; alazba2023deep, ).
@@
In recent years, in the field of Natural Language Processing (NLP), increased computational power, advanced machine learning techniques, and access to very large-scale data have led to the emergence of Large Language Models (LLMs). The exceptional performance of LLMs in NLP tasks (e.g., text classification, generation, and summarization) has garnered significant attention and interest from the Software Engineering (SE) domain. Researchers have begun employed LLMs, particularly code-specific LLMs like CodeBERT-Base (feng2020codebert, ) and CodeT5 (wang2021codet5, ), using fine-tuning techniques to tackle various SE tasks, such as code review (lu2023llamareviewer, ) and vulnerability repair (le2024study, ). Among all the fine-tuning methods, Parameter-Efficient Fine-Tuning (PEFT) has emerged as a promising approach, offering an efficient alternative to full fine-tuning, particularly when computational resources are limited. PEFT methods have shown potential in adapting LLMs to various downstream tasks with fewer parameters and reduced resource requirements, thus expanding the practical applicability of LLMs in diverse SE contexts. Previous studies have explored the efficiency of PEFT methods for Language Models (LMs) on automated program repair (li2024comprehensive, ), code generation (liu2023empirical, ), code search (wang2023one, ), and etc. However, to the best of our knowledge, there is currently no comprehensive research exploring the use of LLMs for code smell detection, nor any in-depth studies analyzing the application of PEFT methods in this area.
@@
Moreover, utilizing PEFT methods for code smell detection can facilitate just-in-time code smell detection. Traditional heuristic-based methods for code smell detection are often applied after the software systems have been completed (designitejava, ; pmd, ; marinescu2005iplasma, ). As a result, code smells are typically discovered at later stages of development, requiring refactoring that involves modifying all relevant parts of the system, thereby increasing the cost and complexity of addressing these code smells. In contrast, incorporating LLMs with PEFT techniques allows for real-time detection of code smells, enabling developers to identify and address potential code smells as they write code. For instance, upon completing a function, developers can immediately receive feedback on whether it contains a code smell. This proactive approach contributes to maintaining high code quality throughout the development life cycle, leading to more efficient and effective code maintenance and improvement (ardimento2021temporal, ; ardimento2021transfer, ).
@@
To this end, we conduct a series of experiments on fine-tuning both small and large LMs with PEFT methods on the code smell detection task. In this paper, we define small LMs as those with fewer than 1B parameters, while LMs with 1B parameters or more are considered large LMs. Specifically, we construct high-quality Java code smell datasets comprising two types of widely studied and sophisticated method-level code smells, i.e., Complex Conditional (CC) and Complex Method (CM), as no existing datasets are available. We then apply various PEFT methods, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$, to fine-tune four popular small LMs and six cutting-edge LLMs with parameter sizes ranging from 125M to 7B. The effectiveness of these PEFT methods is assessed for code smell detection, and their performance is compared to that of full fine-tuning in terms of both effectiveness and peak GPU memory usage. Additionally, we investigate the performance of PEFT techniques under different hyper-parameter settings and evaluate their effectiveness in low-resource scenarios with varying numbers of training samples on the method-level code smell detection task. Finally, we employ a heuristics-based static analysis tool alongside two state-of-the-art DL-based code smell detection tools to compare the performance of PEFT methods with existing approaches.
@@
  • We mine Java repositories from GitHub and employ a traditional heuristic-based detector, along with manual validation, to create high-quality datasets for two types of method-level code smells, i.e., Complex Conditional and Complex Method.
  • We conduct a comprehensive evaluation of multiple PEFT methods on both small and large LMs for code smell detection. Our results indicate that LLMs, when fine-tuned, do not always outperform small LMs. Additionally, PEFT methods achieve better or comparable effectiveness compared to full fine-tuning.
@@
  • We compare a heuristic-based detector and two best-in-class DL-based code smell detection tools with both small and large LMs using PEFT methods on code smell detection. Our results demonstrate that both small and large LMs with PEFT methods consistently outperform the existing heuristic- and DL-based approaches in detecting the method-level code smells, highlighting the superiority of PEFT methods.
  • We provide actionable suggestions and insights for both developers and researchers for using PEFT methods on method-level code smell detection.
@@ 2.3. Parameter-Efficient Fine-Tuning (PEFT) Methods
Instead of full fine-tuning, pre-trained LMs could be tailored to downstream tasks by Parameter-Efficient Fine-Tuning (PEFT) methods. PEFT methods involve optimizing only a subset of model parameters or incorporating external modules for new tasks while keeping the majority of parameters frozen. Compared to full fine-tuning, PEFT methods can reduce the substantial computational and data demands of LLMs and achieve comparable or even better results (han2024parameter, ). The right side of Figure 2 illustrates several state-of-the-art PEFT methods employed in this study, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$.
@@ 3. Methodology
Figure 2. Left: architecture of our proposed method for method-level code smell detection. Right: a detailed illustration of the PEFT module structure for the four PEFT methods.
@@ 3.1. Research Goal
Figure 2 illustrates the architecture of our LLM-based code smell detector using PEFT methods. We frame the method-level code smell detection task as a binary classification problem, where the input is a piece of code. A sequence classification head is added on top of the LLM to handle the classification task. For model fine-tuning, we use four distinct PEFT methods. The model then outputs either 00 or $1$, representing the absence or presence of the specific type of code smell.
@@ 3.2. Research Questions
This study aims to evaluate the effectiveness of state-of-the-art PEFT methods (prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$) in small LMs and LLMs for method-level code smell detection. To achieve this goal, we conducted a set of experiments to answer the following Research Questions (RQs):
RQ1: What is the effectiveness of different PEFT methods and full fine-tuning on various small LMs and LLMs? Previous studies have shown remarkable results using PEFT methods to fine-tune small LMs for software engineering tasks like code clone detection and defect detection (liu2023empirical, ; liu2024delving, ; li2024comprehensive, ; wang2023one, ; weyssow2023exploring, ). However, these studies have not paid attention to the effectiveness of PEFT methods for code smell detection. With this RQ, we aim to compare and validate the effectiveness of PEFT methods across multiple LMs of varying parameter sizes (see Section 3.4) specifically for code smell detection.
RQ2: What is the effectiveness of PEFT methods with different hyper-parameters? When fine-tuning LMs for downstream tasks, the configuration of hyper-parameters plays a crucial role in determining the final results by affecting the number of trainable parameters. In prompt tuning and prefix tuning, the number of virtual tokens serves as a critical hyper-parameter (lester2021power, ; li2021prefix, ). For LoRA, the key hyper-parameter is the rank $r$ (hu2022lora, ), while in $(IA)^{3}$, it is the selection of activated modules (liu2022few, ). Due to our limited computational resources and the popularity of LoRA compared to other PEFT methods (gupta2024beyond, ; zhang2024star, ), we decided to employ LoRA to fine-tune a small LM and an LLM with different $r$ values in a range of {8, 32, 64, 640, 1,280, 2,560} to explore the effect of hyper-parameter settings on PEFT methods for code smell detection.
@@
RQ4: What is the effectiveness of PEFT methods compared to state-of-the-art code smell detection approaches? Existing code smell detectors include traditional heuristic-based and DL-based methods. However, heuristic-based approaches for code smell detection depend on rules and heuristics derived from software metrics (moha2010decor, ; palomba2015mining, ; palomba2016a, ; fokaefs2007jdeodorant, ), which may suffer from several limitations (e.g. different tools may produce inconsistent results, complicating developers’ refactoring decisions (fontana2012automatic, )). As a result, researchers have been exploring the application of DL techniques as alternatives for code smell detection due to the drawbacks of heuristic-based methods (pecorelli2022on, ), and state-of-the-art DL-based code smell detectors have demonstrated strong effectiveness in identifying code smells (alazba2023deep, ; sharma2021code, ). With this RQ, we aim to assess how LMs utilizing PEFT methods compare against heuristic-based code smell detectors and established DL techniques, including general-purpose LLMs, in terms of their ability to detect and classify code smells accurately.
@@ 3.3. Dataset Construction
In this section, we elaborate on the process with the following steps of constructing the high-quality datasets for two types of method-level code smells used to fine-tune both small and large LMs for code smell detection. We created our own datasets because most existing datasets only offer code metrics related to the code smells instead of code snippets, which are not usable for our purposes (sharma2021qscored, ; binh2022mlcodesmell, ; madeyski2023detecting, ). While some code smell datasets provide valid code samples, they do not include Complex Conditional and Complex Method (madeyski2020mlcq, ), two types of method-level code smells. Additionally, the source projects from these existing datasets are quite outdated (fabio2015landfill, ; palomba2018a, ; fontana2016comparing, ; amorim2015experience, ). Considering this, we aim to construct new datasets sourced from the latest trending GitHub projects. Figure 3 illustrates the overview process of dataset construction.

### v2

# A Comprehensive Evaluation of Parameter-Efficient Fine-Tuning on Code Smell Detection

@@ Abstract.
Code smells, which are suboptimal coding practices that can potentially lead to defects or maintenance issues, can negatively impact the quality of software systems. Most existing code smell detection methods rely on heuristics-based or machine learning (ML) and deep learning (DL)-based techniques. However, these techniques have several drawbacks (e.g., unsatisfactory performance). Large Language Models (LLMs) have garnered significant attention in the Software Engineering (SE) field, achieving state-of-the-art performance across a wide range of SE tasks. Parameter-Efficient Fine-Tuning (PEFT) methods, which are commonly used to adapt LLMs to specific tasks with fewer parameters and reduced computational resources, have emerged as a promising approach for enhancing the performance of LLMs in various SE tasks. However, LLMs have not yet been explored for code smell detection, and their effectiveness for this task remains unclear. Furthermore, no comprehensive investigation has been conducted on the efficiency of PEFT methods for code smell detection. In this regard, we systematically evaluate the effectiveness of state-of-the-art PEFT methods on both small and large Language Models (LMs) for code smell detection. To begin, we constructed high-quality Java code smell datasets sourced from GitHub. We then fine-tuned four small LMs (SLMs) and five LLMs using various PEFT techniques, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$, for code smell detection. Our comparison against full fine-tuning revealed that PEFT methods not only achieve comparable or better effectiveness but also consume less peak GPU memory. Our analysis further explored the performance of SLMs versus LLMs in the context of code smell detection. Generally, LLMs outperform SLMs on detecting Complex Conditional, Complex Method, and Feature Envy. However, SLMs demonstrate superior performance on detecting Data Class. We also investigated the impact of varying hyper-parameter settings and the availability of training data. Our results show that increasing the number of trainable parameters does not necessarily improve the effectiveness of PEFT methods for code smell detection. In contrast, we observed significant performance gains when the training dataset size was increased. Finally, our results show that LMs with PEFT methods outperform traditional heuristic-based detectors, advanced DL-based approaches, and general-purpose LLMs under In-Context Learning in the task of code smell detection. Specifically, we observe improvements in MCC ranging from 4.11% to 19.25% across the four types of code smells studied, demonstrating a more effective solution for this task. In conclusion, our work provides a comprehensive evaluation of PEFT methods on both small and large LMs for code smell detection. The results highlight the efficiency and effectiveness of PEFT methods in this domain, offering valuable insights for optimizing LMs for code smell detection. We suggest that the choice of PEFT method for code smell detection should be guided by the model used, data availability, and computational resources. Our replication package is available at (replicationPackage, ), enabling others to reproduce and extend our findings.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: TOSEMjournalvolume: 0journalnumber: 0article: 0publicationmonth: 0ccs: Software and its engineering Software maintenance tools
@@ 1. Introduction
To address these issues, many researchers have incorporated machine learning (ML) and deep learning (DL) algorithms to mitigate the performance instability in code smell detection that arises from the manual threshold-setting process in traditional detectors (fontana2013code, ; fontana2016comparing, ; sharma2021code, ; liu2021deep, ). By learning to discern between smelly and non-smelly code elements, ML and DL techniques offer a solution to the subjectivity problem inherent in heuristics-based methods (nucci2018detecting, ). However, despite these advancements, ML and DL methods still face significant challenges in achieving considerable performance for code smell detection (pereira2022code, ; alazba2023deep, ).
@@
In recent years, in the field of Natural Language Processing (NLP), increased computational power, advanced machine learning techniques, and access to very large-scale data have led to the emergence of Large Language Models (LLMs). The exceptional performance of LLMs in NLP tasks (e.g., text classification, generation, and summarization) has garnered significant attention and interest from the Software Engineering (SE) domain. Researchers have begun employed LLMs, particularly code-specific LLMs like CodeBERT-Base (feng2020codebert, ) and CodeT5 (wang2021codet5, ), to tackle various SE tasks, such as code review (lu2023llamareviewer, ) and vulnerability repair (le2024study, ). However, directly using LLMs without any task-specific fine-tuning for code smell detection yields suboptimal results (mesbah2025leveraging, ; silva2024detecting, ; sadik2025benchmarking, ), as these models lack adaptation to the unique demands of the task.
@@
Fine-tuning enables the adaptation of Language Models (LMs) to specific downstream tasks by updating model parameters, allowing them to better capture task-relevant features and improve performance. Among all the fine-tuning methods, Parameter-Efficient Fine-Tuning (PEFT) has emerged as a promising approach, offering an efficient alternative to full fine-tuning, particularly when computational resources are limited. PEFT methods have shown potential in adapting LLMs to various downstream tasks with fewer parameters and reduced resource requirements, thus expanding the practical applicability of LLMs in diverse SE contexts. Previous studies have explored the efficiency of PEFT methods for LMs on automated program repair (li2024comprehensive, ), code generation (liu2023empirical, ), code search (wang2023one, ), etc. However, to the best of our knowledge, no prior research has focused on improving the performance of LLMs for code smell detection, nor has any in-depth study been conducted analyzing the application of PEFT methods in this area. Additionally, identifying the code smells selected in this study requires both method- and class-level code understanding. While existing research on PEFT methods has predominantly focused on code-related tasks at the method level (li2024comprehensive, ; liu2023empirical, ; wang2023one, ; liu2024delving, ; weyssow2023exploring, ), a significant gap remains in assessing their effectiveness at the class level. By leveraging the task of code smell detection, our study aims to fill this gap by systematically exploring the capacities of LMs with PEFT methods in class-level code understanding. Furthermore, utilizing PEFT methods for code smell detection can facilitate just-in-time code smell detection. Traditional heuristic-based methods for code smell detection are often applied after the software systems have been completed (designitejava, ; pmd, ; marinescu2005iplasma, ). As a result, code smells are typically discovered at later stages of development, requiring refactoring that involves modifying all relevant parts of the system, thereby increasing the cost and complexity of addressing these code smells. In contrast, incorporating LLMs with PEFT techniques allows for real-time detection of code smells, enabling developers to identify and address potential code smells as they write code. For instance, upon completing a function, developers can immediately receive feedback on whether it contains a code smell. This proactive approach contributes to maintaining high code quality throughout the development life cycle, leading to more efficient and effective code maintenance and improvement (ardimento2021temporal, ; ardimento2021transfer, ).
@@
To this end, we conduct a series of experiments on fine-tuning both small and large LMs with PEFT methods on the code smell detection task. In this paper, we define small LMs (SLMs) as those with fewer than 1B parameters, while LMs with 1B parameters or more are considered LLMs. Specifically, we construct high-quality Java code smell datasets comprising four types of widely studied and sophisticated code smells at both the method and class level, i.e., Complex Conditional (CC), Complex Method (CM), Feature Envy (FE), and Data Class (DC). These datasets are developed to address the lack of suitable existing resources, as most available datasets either do not conform to the required data formats or are based on outdated codebases. We then apply various PEFT methods, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$, to fine-tune four popular SLMs and five cutting-edge LLMs with parameter sizes ranging from 125M to 7B. The effectiveness of these PEFT methods is assessed for code smell detection, and their performance is compared to that of full fine-tuning in terms of both effectiveness and peak GPU memory usage. Additionally, we investigate the performance of PEFT techniques under different hyper-parameter settings and evaluate their effectiveness in low-resource scenarios with varying numbers of training samples on the code smell detection task. Finally, we employ a heuristics-based static analysis tool alongside two state-of-the-art DL-based code smell detection tools to compare the performance of PEFT methods with existing approaches.
@@
  • We mine Java repositories from GitHub and employ traditional heuristic-based detectors, along with a two-stage manual review, to create high-quality and balanced datasets for four types of code smells, i.e., Complex Conditional, Complex Method, Feature Envy, and Data Class.
  • We conduct a comprehensive evaluation of multiple PEFT methods on both small and large LMs for code smell detection. Our results indicate that LLMs, when fine-tuned, do not always outperform SLMs. Moreover, PEFT methods achieve better or comparable effectiveness compared to full fine-tuning. Overall, LMs with PEFT methods exhibit promising capabilities in method- and class-level code understanding.
@@
  • We compare a heuristic-based detector, two best-in-class DL-based detection tools, and two general-purpose LLMs using In-Context Learning (ICL) against our PEFT-based method on code smell detection. Our results demonstrate that LMs with PEFT methods consistently outperform all baselines across multiple code smell types, highlighting the superiority of PEFT methods.
  • We provide actionable suggestions and insights for both developers and researchers for using PEFT methods on code smell detection.
@@ 2.3. Parameter-Efficient Fine-Tuning (PEFT) Methods
Instead of full fine-tuning, pre-trained LMs could be tailored to downstream tasks by Parameter-Efficient Fine-Tuning (PEFT) methods. PEFT methods involve optimizing only a subset of model parameters or incorporating external modules for new tasks while keeping the majority of parameters frozen. Compared to full fine-tuning, PEFT methods can reduce the substantial computational and data demands of LLMs and achieve comparable or even better results (han2024parameter, ).
The right side of Figure 2 illustrates several state-of-the-art PEFT methods employed in this study, including prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$. These methods were selected based on several considerations. First, they are all supported by the widely used Hugging Face PEFT library [^1] https://github.com/huggingface/peft, which provides a standardized and efficient framework for applying PEFT to LMs. Second, according to the PEFT library’s documentation [^2] https://huggingface.co/docs/peft/index, PEFT methods are categorized into three main types: Prompt-based methods, LoRA methods, and $(IA)^{3}$. The four selected methods span these categories, making them representative of the major approaches currently available. Finally, these methods are not only recent but also widely adopted in contemporary research (weyssow2023exploring, ; li2024comprehensive, ; mannisto2025comparative, ; zhou2024empirical, ; taylor2022galactica, ; esmaeili2024empirical, ; wistuba2024choice, ). Their demonstrated effectiveness across various tasks and their popularity in both the NLP and SE communities make them ideal candidates for this study. By selecting these four methods, we ensure that our experiments cover a range of the most promising and commonly adopted techniques in PEFT.
@@ 3. Methodology
Figure 2. Left: architecture of our proposed method for code smell detection. Right: a detailed illustration of the PEFT module structure for the four PEFT methods.
@@ 3.1. Research Goal
Figure 2 illustrates the architecture of our LLM-based code smell detector using PEFT methods. We frame the code smell detection task as a binary classification problem, where the input is a piece of code. A sequence classification head is added on top of the LLM to handle the classification task. For model fine-tuning, we use four distinct PEFT methods. The model then outputs either 00 or $1$, representing the absence or presence of the specific type of code smell.
@@ 3.2. Research Questions
This study aims to evaluate the effectiveness of state-of-the-art PEFT methods (prompt tuning, prefix tuning, LoRA, and $(IA)^{3}$) in SLMs and LLMs for code smell detection. To achieve this goal, we conducted a set of experiments to answer the following Research Questions (RQs):
RQ1: What is the effectiveness of different PEFT methods and full fine-tuning on various SLMs and LLMs? Previous studies have shown remarkable results using PEFT methods to fine-tune SLMs for software engineering tasks like code clone detection and defect detection (liu2023empirical, ; liu2024delving, ; li2024comprehensive, ; wang2023one, ; weyssow2023exploring, ). However, these studies have not paid attention to the effectiveness of PEFT methods for code smell detection. With this RQ, we aim to compare and validate the effectiveness of PEFT methods across multiple LMs of varying parameter sizes (see Section 3.4) specifically for code smell detection.
RQ2: What is the effectiveness of PEFT methods with different hyper-parameters? When fine-tuning LMs for downstream tasks, the configuration of hyper-parameters plays a crucial role in determining the final results by affecting the number of trainable parameters. In prompt tuning and prefix tuning, the number of virtual tokens serves as a critical hyper-parameter (lester2021power, ; li2021prefix, ). For LoRA, the key hyper-parameter is the rank $r$ (hu2022lora, ), while in $(IA)^{3}$, it is the selection of activated modules (liu2022few, ). Due to our limited computational resources and the popularity of LoRA compared to other PEFT methods (gupta2024beyond, ; zhang2024star, ), we decided to employ LoRA to fine-tune a SLM and an LLM with different $r$ values in a range of {8, 32, 64, 640, 1,280, 2,560} to explore the effect of hyper-parameter settings on PEFT methods for code smell detection.
@@
RQ4: What is the effectiveness of PEFT methods compared to state-of-the-art code smell detection approaches? Existing code smell detectors include traditional heuristic-based and DL-based methods. Additionally, the use of general-purpose LLMs with In-Context Learning (ICL) has become increasingly popular in SE tasks (zhou2024large, ; shin2023prompt, ; li2025large, ; xu2024unilog, ), and such models can also be leveraged for code smell detection. With this RQ, we aim to assess how LMs utilizing PEFT methods compare against heuristic-based code smell detectors, established DL techniques, and general-purpose LLMs under zero- and few-shot settings for the task of code smell detection.
@@ 3.3. Dataset Construction
In this section, we elaborate on the process with the following steps of constructing the high-quality and balanced datasets for four types of code smells used to fine-tune both small and large LMs for code smell detection. We created our own datasets because most existing datasets only offer code metrics related to the code smells instead of code snippets, which are not usable for our purposes (sharma2021qscored, ; binh2022mlcodesmell, ; madeyski2023detecting, ). While some code smell datasets provide valid code samples, they do not cover the specific types of code smells we focus on (e.g., Complex Conditional and Complex Method) (madeyski2020mlcq, ). To better align with current software development practices, we constructed new datasets based on trending GitHub projects created within the past six years. Furthermore, since data imbalance may negatively impact model performance on code smell detection (khleel2024improving, ; nandini2025improving, ; fontana2016comparing, ; nunes2024tuning, ), we ensured that our constructed datasets contain an equal number of positive and negative samples. Figure 3 illustrates the overview of the dataset construction process.